signmentsy/355039 Kimmel, Financial Accounting, 8e E xercise 6-11 (Part Level Submission) Suppose this information is available for PepsiCo, Inc, for 2015, 2016, and 2017 (in Beginning inventory Ending inventory Cost of goods sold Sales revenue $2,000 $2,400 $2,600 2,400 2,600 2,0o 17,738 20,209 20,394 39,307 43,073 43,959 ▼ (al) Your answer s incorrect. Try again. Cakculate the inventory turnover for Pepsico, Iinc for 2015, 2016, and 2017. (Round inventory turnover to 1 decimal place, e.q. S.t.) Inventory turnover Attempts: 101 S used sav a·LATER (a3) The parts of thes question must be completed order. Thes part wil be available when you complete the part above (a3) The parts of this question must be completed in order. Thes part will be available when you compiete the art above
Answer #1

Inventory turnover = Cost of Goods Sold / Average inventory

2015:

Inventory turnover = $17,738 / [($2,000 + $2,400) / 2] = 8.1

2016:

Inventory turnover = $20,209 / [($2,400 + $2,600) / 2] = 8.1

2017:

Inventory turnover = $20,394 / [($2,600 + $2,700) / 2] = 7.7